OpenStudy (anonymous):

Solve the equation for y and then graph it. y - 5 = 2x +5 +5 _____________ y=2x+5; how do you graph it?

OpenStudy (anonymous):

You know that it hits the y axis at 5.

OpenStudy (anonymous):

You know that it rises 2 blocks for every one block.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but does it rise to the right or left?

OpenStudy (anonymous):

So make a point at \((0,5)\) and another point at \((1,7)\) and perhaps another at \((-1,3)\). Then use a ruler to do the rest.

OpenStudy (anonymous):

The easiest way is to just let \(x=1\) and see what \(y\) becomes. In this case:\[ y = 2x+5=2(1)+5 = 2+5=7 \]So \(x=1,y=7\) means the line hits the point \((1,7)\)